OpenStudy (roku):

quick question: work out the following and tell me the answer you get. here goes A flask contains 0.750 mole of CO2. what mass of carbon dioxide is in the sample.

OpenStudy (roku):

question number two: how many moles of methane are there in a 23.6 gram sample

sam (.sam.):

What do you get for #1?

OpenStudy (anonymous):

No : of moles = Given mass/ Molar mass 0.75 = x/44 x = 33 gm

OpenStudy (anonymous):

Molar Mass for C: 12, O: 16, H:1 Mass of CO2 = 0.75*(12+2(16) x = 33 g for Methane (CH4), Mol CH4 = 23.6/ (12+4) = 1. 475 mol

OpenStudy (roku):

got the same answers which means i understand somewhat posting another question please respond

OpenStudy (anonymous):

It's not that hard (: my chem teacher gave us this diagram, it makes it simpler to understand : So basically to get the mole, you have to divided the given info (mass, volume, no. of particles) with the molar mass, temperature, or avogrado's number (: